### v1

# VecCity: A Taxonomy-guided Library for Map Entity Representation Learning [Experiment, Analysis & Benchmark]

Abstract.
Electronic maps consist of diverse entities, such as points of interest (POIs), road segments, and land parcels, playing a vital role in applications like ITS and LBS. Map entity representation learning (MapRL) generates versatile and reusable data representations, providing essential tools for efficiently managing and utilizing map entity data. Despite the progress in MapRL, two key challenges constrain further development. First, existing research is fragmented, with models classified by the type of map entity, limiting the reusability of techniques across different tasks. Second, the lack of unified benchmarks makes systematic evaluation and comparison of models difficult. To address these challenges, we propose a novel taxonomy for MapRL that organizes models based on functional modules—such as encoders, pre-training tasks, and downstream tasks—rather than by entity type. Building on this taxonomy, we present a taxonomy-driven library, VecCity, which offers easy-to-use interfaces for encoding, pre-training, fine-tuning, and evaluation. The library integrates datasets from nine cities and reproduces 21 mainstream MapRL models, establishing the first standardized benchmarks for the field. VecCity also allows users to modify and extend models through modular components, facilitating seamless experimentation. Our comprehensive experiments cover multiple types of map entities and evaluate 21 VecCity pre-built models across various downstream tasks. Experimental results demonstrate the effectiveness of VecCity in streamlining model development and provide insights into the impact of various components on performance. By promoting modular design and reusability, VecCity offers a unified framework to advance research and innovation in MapRL. The code is available at https://github.com/Bigscity-VecCity/VecCity.
1. Introduction
In the era of mobile internet, electronic maps have become a foundational platform for a wide range of applications, including intelligent transportation systems and location-based services. An electronic map consists of map entities, such as points of interest (POIs), road segments, and land parcels. These entities encapsulate complex geospatial data, spatial relationships, and geometric-topological structures, presenting significant challenges for data representation. The effective representation of map entities has become a critical and enduring research focus within spatiotemporal data analysis and geographic information systems (GIS).
In the literature, early studies represented map entities as records in relational databases or as files containing geographic information attributes, such as PostGIS (Team, 2023), Shapefiles (Environmental Systems Research Institute, 1998), and GeoJSON (Butler et al., 2016). With this representation approach, application services relying on map entities are required to develop task-specific models or algorithms to process the entities’ attributes and extract geographic relationships within or between them. However, the task-specific nature of these data processing models restricts their generalizability and reusability across different applications and scenarios.
Figure 1. ArXiv submissions on MapRL over time.
In recent years, pre-trained representation learning has emerged as a powerful approach for generating versatile, task-independent data representations. It has achieved remarkable success across various domains, including natural language processing (NLP) (Radford et al., 2018; Mikolov et al., 2013; Peters et al., 2018; Pennington et al., 2014) and computer vision (CV) (Chen et al., 2020; He et al., 2020, 2016; Dosovitskiy et al., 2021; Vincent et al., 2010). In the field of spatiotemporal data analysis, a growing number of studies have embraced pre-trained representation learning methods to construct generic representations of map entities (Mai et al., 2022; Chen et al., 2024a; Zhang et al., 2024; Li et al., 2024b; Ning et al., 2024). These efforts have given rise to an emerging research area known as Map Entity Representation Learning (MapRL). As shown in Fig. 1, the number of MapRL papers has been steadily increasing, reflecting the growing interest and potential impact of this field. However, despite these rapid advancements, several structural challenges require further exploration to unlock the potential of MapRL.
Challenge 1: Fragmented Research Fields. Electronic maps consist of various types of map entities, such as POIs (points), road segments (polylines), and land parcels (polygons). Existing research often treats the modeling of these entities as separate areas, limiting the reusability of techniques, methods, and modules across fields. However, real-world electronic map applications typically require the integrated use of all three entity types. This segregation not only creates obstacles for practical applications but also forms implicit academic exchange barriers between subfields, thereby constraining the progress of MapRL research.
Challenge 2: Lack of Standardized Benchmark. Although numerous MapRL models have been proposed in the literature, their performance is often evaluated on different datasets and under varying experimental settings. Unlike the CV and NLP fields, this domain lacks standardized datasets and benchmarks. This absence hinders fair performance comparisons between models and limits the ability of researchers to establish guiding design principles through consistent and rigorous model evaluation.
To address the first challenge, we propose a novel taxonomy for MapRL models, namely method-based taxonomy, which organizes the essential components of an MapRL model into four key elements: Map Data, Encoder Models, Pre-training Tasks, and Downstream Tasks. The map data is further divided into Map Entity Data and Auxiliary Data. Encoder models are categorized into three types: Token-based, Graph-based, and Sequence-based models, each supporting different pre-training tasks. A key insight of our taxonomy is that the two core elements—Encoder Models and Pre-training Tasks—are not tightly coupled with specific map entity types. Only the downstream tasks depend on the type of map entity. The method-based taxonomy represents a significant departure from the traditional classification of MapRL models, i.e., type-based taxonomy, which primarily focuses on map entity types.
To address the second challenge, we propose an MapRL model development toolkit based on our taxonomy, namely the VecCity library. The library organizes map entities and auxiliary data (i.e., the map data element) into three types of atomic files. Using these atomic files, we format data from nine cities into a unified structure. the VecCity library further provides four interface functions to standardize the implementation processes (i.e., the encoder model, pre-training task, and downstream task elements) of MapRL models. Leveraging these interfaces, we have implemented 21 mainstream MapRL models. Additionally, we conduct a comprehensive evaluation of these models across the datasets from the nine cities, offering valuable insights to guide future research.
The main contributions of this paper are as follows:
  • Novel Taxonomy. We present a novel taxonomy of MapRL models that goes beyond the traditional classification based on map entities, offering a unified classification system applicable to MapRL models across various types of map entities.
  • Easy-to-use Toolkit Library. Guided by our taxonomy, we propose a toolkit, VecCity, which unifies the development process of MapRL models. To the best of our knowledge, VecCity is the first unified model toolkit specifically designed for MapRL.
  • Standard Benchmark. Using VecCity, we implement 21 mainstream MapRL models and evaluate their performance on datasets from nine cities. This work establishes the first standard benchmark for all types of MapRL models and provides valuable insights to guide future research.
Figure 2. The framework of a MapRL model.
2. Taxonomy of MapRL
In this section, we present the proposed method-based taxonomy. Figure 2 outlines the general framework for MapRL. The training process of a typical MapRL model consists of two key stages: pre-training and fine-tuning. During pre-training, the MapRL model takes map data as input, using an encoder model to transform them into representation vectors. The encoder’s parameters are optimized through pre-training tasks. In the fine-tuning stage, downstream tasks are employed to further refine the encoder. This pipeline defines four core components of an MapRL model: Map Data, Encoder Models, Pre-training Tasks, and Downstream Tasks.
2.1. Map Data
The map data consists of map entities and auxiliary data.
2.1.1. Map Entities
As illustrated in Fig. 3, an electronic map consists of three geometric elements: Points, Polylines, and Polygons, each representing a distinct type of map entity.
  • Points / POI. A point element represents a Point of Interest (POI), identified by a location coordinate, i.e., $(latitude,longitude)$.
  • Polylines / Road Segment. A polyline element represents a Road Segment, composed of a sequence of connected line segments, with each endpoint defined by location coordinates.
  • Polygons / Land Parcel. A polygon element represents a Land Parcel, defined by a set of connected line segments forming a closed boundary.
For each map entity, a set of features describes its properties. For instance, the category feature of a POI differentiates the services it offers, such as scenic spots, transit stations, or shopping facilities. Road segments are characterized by features like speed limits, lane counts, road grade, etc. Based on these definitions, we formally define electronic maps and map entities as follows.
Definition 0 (Electronic Maps and Map Entities).
An electronic map is defined as a set of map entities, $V=\{v_{1},\ldots,v_{i},\ldots,v_{I}\}$, where each entity $v_{i}$ consists of four components: ID, type, geometric shape, and feature. The ID serves as the unique identifier of $v_{i}$. The type specifies the entity’s geometric form – point, polyline, or polygon. The geometric shape provides the coordinates that define its structure, while the feature captures its various attributes.
Figure 3. Urban map with different map entities.
2.1.2. Auxiliary Data
Besides map entities, two types of auxiliary data are commonly used in MapRL models: Trajectories and Relation Networks.
Definition 0 (Trajectories).
Trajectories refer to the mobility records of humans or vehicles. A trajectory is defined as a sequence of locations with corresponding timestamps, i.e., $tr=(<l_{tr_{1}},t_{tr_{1}}>,\ldots,<l_{tr_{k}},t_{tr_{k}}>,\ldots,<l_{tr_{K}},t_{tr_{K}}>)$, where $l_{tr_{k}}$ and $t_{tr_{k}}$ are the location and timestamp for the $k$-th trajectory sample.
Trajectory data can be divided into two classes based on the type of location samples $l_{tr_{k}}$. $i$) Check-in Trajectories: These use POIs as samples and are collected through social networks or location-based service (LBS) apps. They log a user’s location only upon check-in, resulting in a low sampling rate, often requiring days to capture a single point. MapRL models leverage these trajectories to infer users’ personal preferences. $ii$) Coordinate Trajectories: These rely on latitude and longitude coordinates as samples. Collected by intelligent transportation systems (ITS) through GPS devices or mobile phones, they provide high-frequency sampling along with real-time speed and direction data. These trajectories are typically aligned with road segments using map matching algorithms (Yang and Gidofalvi, 2018) and facilitate route reconstruction and dynamic state estimation for road segments.
Definition 0 (Relation Network).
A relation network represents map entities as a directed graph $G=\{V,\bm{E}\}$, where $V=\{v_{1},\ldots,v_{I}\}$ is the set of vertices corresponding to map entities. $\bm{E}\in\mathbb{R}^{I\times I}$ is an adjacency matrix, with each element $e_{ij}$ indicating the relationship from $v_{i}$ to $v_{j}$.
Network relations can be categorized into two types: $i$) Geographical Relations (GR): These leverage spatial relationships between map entities to construct the adjacency matrix. Examples include connections between road segments in a road network (Zhang and Long, 2023) and distances between POIs (Zhao et al., 2017; Du et al., 2019). $ii$) Social Relations (SR): These capture user behaviors, such as mobility trajectories, to form adjacency matrices. A common example is Origin-Destination (OD) flows, widely used in MapRL models (Zhang et al., 2020; Zhou et al., 2023), where the number of trajectories originating from $v_{i}$ and terminating at $v_{j}$ defines the value of $e_{ij}$ in the adjacency matrix.
2.2. Encoder Models and Pre-training Tasks
Encoder models and pre-training tasks are two core components of a pre-trained MapRL model, responsible for converting input map data into representation vectors for map entities. Our taxonomy classifies encoder models into three types: token-based, graph-based, and sequence-based, designed to model features, relation networks, and trajectory sequences of map entities, respectively. Each encoder type corresponds to different pre-training tasks tailored to its structure and purpose.
2.2.1. Token-based Models
Encoders. The token-based encoder generates representation vectors directly from the features of map entities. For discrete features, it applies one-hot encoding, and for continuous features, the value range is divided into consecutive bins, each represented by a one-hot code. Given $K$ features for a map entity, let the one-hot code of the $k$-th feature be $\bm{f}_{k}\in\{0,1\}^{F_{k}}$ (where $F_{k}$ is the dimensionality). The encoder employs a learnable embedding matrix $\bm{R}_{k}\in\mathbb{R}^{F_{k}\times d}$ for each feature to generate the entity’s representation:
(1) $\small\bm{r}=\left\|{}_{k=1}^{K}\bm{f}_{k}^{\top}\cdot\bm{R}_{k}\right.,$
where $\|_{k=1}^{K}$ denotes the concatenation of all features’ embeddings.
In the literature, the novelty of existing token-based methods lies in how effectively they extract valuable features from map data. These features can be categorized into three types. $i$) Spatial Features: These capture the locations and spatial structures of map entities. Spatial coordinates are the most fundamental spatial features. Hier (Shimizu et al., 2020) incorporates spatial structures by embedding hierarchical spatial grids into the representations of rasterized maps. $ii$) Temporal Features: A map entity’s characteristics can vary at different times of the day. For example, a restaurant exhibits different behaviors during working hours versus lunchtime. To capture these variations, many token-based encoders divide the day into segments (e.g., hourly or half-hourly) and assign each segment a unique embedding to represent temporal features (Wan et al., 2021; Zhao et al., 2017; Gong et al., 2023; Ma et al., 2024; Liu et al., 2016). $iii$) Semantic Features: These reflect the intrinsic properties of map entities, such as the functional types of POIs or the road types of segments. Some land parcel MapRL methods further enhance semantic features by counting the number of POI categories within a parcel (Zhang et al., 2022).
Pre-training Tasks. For token-based encoders, the corresponding pre-training tasks include three types:
  • Token Relation Inference (TokRI). This task aims to predict the relationship between two map entities based on their representation vectors. Specifically, given the representations $\bm{r}_{i}$ and $\bm{r}_{j}$ of two entities, the TokRI task predicts their relationship as $\hat{y}_{ij}=\phi(\bm{r}_{i},\bm{r}_{j})$, where $\phi(\cdot)$ is the prediction function, and $\hat{y}_{ij}\in[0,1]$ is the predicted probability of a relation existing. The encoder parameters are optimized during pre-training using cross-entropy loss.
  • Token Relation Contrastive Learning (TRCL). This task ensures that entities with closer relationships are more likely to have similar representations. Given a representation sample $\bm{c}_{i}$, another sample $\bm{c}_{j}$ is treated as a positive sample if the corresponding entities $v_{i}$ and $v_{j}$ share a close relation; otherwise, $\bm{c}_{j}$ is treated as a negative sample. The encoder parameters are optimized using a contrastive learning loss, such as the InfoNCE loss (Oord et al., 2018).
  • Augmented Token Contrastive Learning (AToCL). These tasks use data augmentation techniques to generate different views of a given entity’s representation. A contrastive learning loss is then applied to maximize the similarity between views of the same entity while minimizing the similarity between different samples.
For TokRI and TRCL tasks, the novelty of existing models stems from their methods to constructing token relations. The novelty of existing TokRI and TRCL tasks lies in their approaches to constructing token relations, which can be categorized into two methods: $i$) Geographical Distance-Based Relations: This method uses spatial proximity to establish token relations. In TokRI tasks, RN2Vec (Wang et al., 2020) defines a relationship between entities if their geographical distance is below a preset threshold. In TRCL tasks, Teaser (Zhao et al., 2017) selects the $K$ nearest entities as positive samples for contrastive learning. $ii$) Auxiliary Data-Based Relations: This method leverages additional data, such as trajectories, to define relations. In TokRI tasks, P2Vec (Feng et al., 2017) and Tale (Wan et al., 2021) treat a trajectory as a sentence and the entities within it as words, using the Continuous Bag of Words (CBOW) model (Mikolov et al., 2013) to learn representations for “words” (map entities). In TRCL tasks, Teaser (Zhao et al., 2017) employs the SkipGram model (Mikolov et al., 2013) with negative sampling to capture contextual relations along trajectories. When trajectory data is unavailable, “virtual trajectories” can be generated via random walks on relation networks (Wang and Li, 2017). Beyond trajectories, ZEMob (Yao et al., 2018) and DeepMove (Zhou and Huang, 2018) leverage OD flows between entities to identify relationships.
For AToCL tasks, the novelty of existing models lies in their methods for generating augmented samples. For example, ReMVC (Zhang et al., 2022) and ReDCL (Li et al., 2023) augment data by replacing or dropping parts of the anchor entity’s features.
2.2.2. Graph-based Models
Encoders. Graph-based encoders leverage auxiliary data in the form of relation networks (see Definition 3) to capture relationships among map entities. Graph-based encoders typically use the representations generated by token-based encoders as inputs and apply a graph neural network (GNN) to produce refined representation vectors. Specifically, given the map entity set $V=\{v_{1},\ldots,v_{I}\}$, the embeddings generated by the token-based encoder are denoted as $\bm{R}=\{\bm{c}_{1},\ldots,\bm{c}_{I}\}$. The graph-based encoder then takes $\bm{R}$ along with the relation network $G$ to generate the final representations:
(2) $\bm{H}=\left\{\bm{h}_{1},\ldots,\bm{h}_{i},\ldots,\bm{h}_{I}\right\}=\mathrm{GNN}\left(\bm{R},G\right),$
where $\bm{h}_{i}$ is the representation vector for the $i$-th map entity, and $\mathrm{GNN}(\cdot)$ is the GNN-based encoder model. Compared to token-based models, graph-based encoders incorporate the relation networks into the entity representations, enriching their expressiveness.
Existing research on graph-based encoders focuses on developing GNNs capable of capturing complex relationships among map entities. Early studies applied standard GNN models, such as graph convolutional networks (Yu et al., 2018; Wu et al., 2019) and graph attention networks (Song et al., 2022; Zhang et al., 2019b), to explore graphical (Chang et al., 2023; Jiang et al., 2023; Jiao et al., 2020) and social relations (Zhang et al., 2020; Liu et al., 2020). Given the diverse relationships among map entities, some studies adopt more compound GNNs to enhance graph-based encoders. For example, the HREP model constructs a heterogeneous graph with four types of relations to represent land parcels (Zhou et al., 2023). The MGFN model utilizes human mobility patterns to build a multi-graph, assigning a dedicated GNN to each subgraph (Wu et al., 2022). HRNR (Wu et al., 2020) adopts a hierarchical GNN to capture the connection, structure, and function relations of road segments (Wu et al., 2020). Recent advancements also explore relations across different types of map entities. For instance, HRoad employs hypergraph neural networks to capture the relationships between road segments and land parcels (Zhang and Long, 2023).
Pre-training Tasks. For graph-based models, the corresponding pre-training tasks include three types:
  • Node Feature Inference (NFI). The goal of this task is to infer the features of map entities based on the representations extracted by graph-based encoders. Let $\bm{h}_{i}$ denote the representation of the $i$-th map entity. The NFI task predicts the raw features of the corresponding entity as: $\hat{\bm{f}}_{i}=\phi(\bm{h}_{i})$, where $\hat{\bm{f}}_{i}$ is the predicted feature, and $\phi(\cdot)$ is the prediction function, typically implemented as a multi-layer perceptron (MLP). The learning objective for the NFI task is to minimize the prediction error.
  • Graph Autoencoder (GAu). This task employs an autoencoder to reconstruct the adjacency matrix of relation networks. Given the representation vectors of all map entities, denoted as $\bm{H}$, the GAu task reconstructs the adjacency matrix $\bm{E}$ as $\bm{E}$ as $\hat{\bm{E}}=\phi\left(\bm{H}\bm{H}^{\top}\right)$, where $\hat{\bm{E}}$ is the reconstructed adjacency matrix, and $\phi(\cdot)$ is a neural network acting as the decoder. Here, the graph-based encoder serves as the encoder of the autoencoder. The objective is to minimize the difference between $\hat{\bm{E}}$ and $\bm{E}$.
  • Neighborhood Contrastive Learning (NCL). The intuition behind this task is that connected map entities in a relation network should exhibit similar representations. This task treats connected entities as positive samples and unconnected entities as negative samples. Contrastive learning loss functions are then employed to optimize the encoder parameters.
  • Augmented Graph Contrastive Learning (AGCL). AGCL applies data augmentation-based contrastive learning on map entity relation networks. Given a relation network $G$ and its augmented version $\tilde{G}$, the task treats the representations of the same node across these versions as positive samples, while representations of different nodes are considered negative samples.
In the NFI tasks, most of methods use raw entity features as the supervised labels (Zhang and Long, 2023; Wu et al., 2020; Chen et al., 2024c). In particular, GMEL (Liu et al., 2020) converts the OD flows relation network as in/out-flows features for prediction. In the GAu task, the pre-training method design is closely connected to the structure of GNN in the encoder. For example, HRNR (Wu et al., 2020) uses a three-layer hierarchical GNN, requiring different adjacency matrix reconstruction tasks for each layer. HRoad (Zhang and Long, 2023) adopts a hypergraph in its encoder, so its pre-training step incorporates a hyperedge reconstruction task. In NCL tasks, the primary differences between models lie in the relation networks they utilize. For instance, HREP (Zhou et al., 2023) and ReDCL (Li et al., 2023) connect the geographically nearest entities to incorporate spatial information. In AGCL tasks, the choice of data augmentation is critical. Randomly deleting edges can disrupt essential connections. To mitigate this, some methods use edge weights to determine deletion probabilities (Chang et al., 2023; Zhou et al., 2024).
Table 1. Summary of existing MapRL models. "*" indicates the reproduced models in the VecCity library.
Methods Encoders Pretraining Tasks Map Entities Features Aux. Data
Token Graph Seq TokRI TRCL AToCL NFI GAu NCL AGCL TrajP MTR ATrCL POI RS LP Sem Spa Temp Traj Rel
Point of Interest CWAP (Liu et al., 2016) CIT
CAPE (Chang et al., 2018) CIT
SkipG* (Mikolov et al., 2013) CIT
Tale* (Wan et al., 2021) CIT
Teaser* (Zhao et al., 2017) CIT GR
Hier* (Shimizu et al., 2020) CIT GR
P2Vec* (Feng et al., 2017) CIT GR
CACSR* (Gong et al., 2023) CIT
CTLE* (Lin et al., 2021) CIT
Road Segment TrajRNE (Schestakov et al., 2023) GR/SR
DyToast (Chen et al., 2024b) CDT GR
JGRM (Ma et al., 2024) CDT GR
RN2Vec* (Wang et al., 2020) GR
HRNR* (Wu et al., 2020) GR
SARN* (Chang et al., 2023) GR
Toast* (Chen et al., 2021) CDT GR
HRoad* (Zhang and Long, 2023) GR
START* (Jiang et al., 2023) CDT GR
JCLRNT* (Mao et al., 2022) CDT GR
Land Parcel CGAL (Zhang et al., 2019a) GR/SR
DLCL (Du et al., 2019) GR/SR
HUGAT (Kim and Yoon, 2022) GR/SR
Re2Vec (Luo et al., 2022) GR/SR
ReDCL (Li et al., 2023) GR
HAFus (Sun et al., 2024) GR/SR
ReCP (Li et al., 2024a) SR
ZEMob* (Yao et al., 2018) SR
GMEL* (Liu et al., 2020) GR/SR
MGFN* (Wu et al., 2022) SR
HDGE* (Wang and Li, 2017) GR/SR
MVURE* (Zhang et al., 2020) SR
ReMVC* (Zhang et al., 2022) SR
HREP* (Zhou et al., 2023) GR/SR
2.2.3. Sequence-based Models
Encoders. Sequence-based encoders aim to capture temporal dependencies among map entities within trajectory-type auxiliary data. They use sequential deep learning models, such as LSTMs (Sak et al., 2014), GRUs (Cho, 2014), or Transformers (Vaswani et al., 2017), as backbones to encode the embeddings of map entities along the same trajectory into a sequence of representation vectors. These input embeddings are typically generated by token-based or graph-based encoders. Specifically, for map entities along a trajectory $tr$, let their embedding sequence be $\bm{R}_{tr}=(\bm{r}_{tr_{1}},\ldots,\bm{r}_{tr_{k}},\ldots,\bm{r}_{tr_{K}})$. The sequence-based model generates representation vectors for these entities as:
(3) $\small\bm{S}_{tr}=(\bm{s}_{tr_{1}},\ldots,\bm{s}_{tr_{k}},\ldots,\bm{s}_{tr_{K}})=\mathrm{SEQ}(\bm{R}_{tr}),$
where $\mathrm{SEQ}(\cdot)$ is the sequence-based encoder, and $\bm{s}_{tr_{k}}$ is the representation vector for the $k$-th entity along the trajectory $tr$.
In trajectory data, each location sample is typically associated with a timestamp. Integrating temporal information from timestamps into entity representations has become a key focus of sequence-based encoders. Early methods divide time into discrete slots, such as the hour of the day or the day of the week, and use these slot indices to generate input embeddings using token-based encoders (Mao et al., 2022; Feng et al., 2017; Zhao et al., 2017). More advanced models embed timestamps within the components of sequential models. For example, JCLRNT (Mao et al., 2022), Toast (Chen et al., 2021), and CTLE (Lin et al., 2021) incorporate timestamps into the positional encoding of Transformers. Similarly, START (Jiang et al., 2023) integrates the travel time of trajectory through a road segment into the attention of Transformers.
Pre-training Tasks. For sequence-based models, the corresponding pre-training tasks include three types:
  • Trajectory Prediction (TrajP). The goal of this task is to predict the second half of a trajectory using the first half. Given a trajectory $tr$, this task is formally defined as $\hat{tr}_{[k+1:K]}=Pre(\bm{S}_{[1:k]}),~{}\bm{S}_{[1:k]}=Enc(tr_{[1:k]})$, where $Enc(\cdot)$ is the sequence-based encoder, and $Pre(\cdot)$ is the prediction function. Here, $tr_{[1:k]}$ represents the first $k$ samples of the trajectory, and $\bm{S}_{[1:k]}$ is their corresponding sequence of representation vectors. The objective is to minimize the distance between $\hat{tr}_{[k+1:K]}$ and the ground truth.
  • Masked Trajectory Recovery (MTR). Inspired by the masked language model (MLM) used in pre-trained language models (Devlin et al., 2019; Liu et al., 2019), this task randomly masks part of the samples in a trajectory and employs an autoencoder to recover them. Given a trajectory $tr$ and its masked version $\tilde{tr}$, the recovery process is defined as $\hat{tr}=Dec(\tilde{\bm{S}}),~{}\tilde{\bm{S}}=Enc(\tilde{tr})$, where $Enc(\cdot)$ is the sequence-based encoder, and $Dec(\cdot)$ is the decoder, implemented using a deep model. The recovered trajectory is denoted as $\hat{tr}$. The objective is to minimize the distance between $tr$ and $\hat{tr}$.
  • Augmented Trajectory Contrastive Learning (ATrCL). This task creates multiple augmented versions of an original trajectory using data augmentation techniques (Fang et al., 2020; Zhao et al., 2019). Sequence-based encoders are then used to generate representations for both the original and augmented trajectories. A contrastive learning loss is applied to increase the similarity between representations derived from the same original trajectory while decreasing the similarity between those from different trajectories.
In TrajP tasks, the prediction labels can be various. For example, Hier (Shimizu et al., 2020) uses entity IDs as prediction labels, while CTLE (Lin et al., 2021) employs visiting time as labels to incorporate temporal information. In MTR tasks, the main difference across methods lies in how masking positions are chosen. A basic approach is to randomly mask positions within trajectories (Lin et al., 2021), but this overlooks the sequential dependencies between samples. To address this, START (Jiang et al., 2023) and Toast (Chen et al., 2021) mask continuous sub-sequences within a trajectory. In ATrCL tasks, different methods adopt distinct augmentation techniques to enhance trajectory representations. JCLRNT (Mao et al., 2022) generates augmented trajectories by randomly deleting or replacing samples within the trajectory. Considering contextual information, START (Jiang et al., 2023) masks a continuous sub-sequence or replaces it with a new one. CACSR (Gong et al., 2023) introduces white noise into the representations to augment trajectories.
2.2.4. Module Integration of Encoders and Pre-training Tasks
The encoders and pre-training tasks introduced in the previous sections form the fundamental modules for MapRL models. A complete MapRL model can integrate multiple types of encoders and pre-training tasks, leveraging their complementary strengths to enhance representation learning. For the encoder modules, existing models typically follow a “token $\rightarrow$ graph $\rightarrow$ sequence” modeling pipeline. This pipeline first extracts basic features using token-based encoders, then captures relational structures through graph-based encoders, and finally models temporal dependencies with sequence-based encoders. For encoders at each stage of the pipeline, the MapRL model can simultaneously employ multiple encoders for different pre-training tasks. Table 1 summarizes the combinations of encoders and pre-training tasks used in mainstream models.
For the three-stage encoder pipeline, MapRL models follow two pre-training paradigms. $i$) Sequential Training: The encoders are trained in stages – starting with token-based, followed by graph-based, and finally sequence-based – each refining the previous stage’s output. $ii$) Joint Training: All encoders are optimized together using multi-task learning, capturing features, relations, and temporal patterns in an integrated manner. These paradigms offer flexibility in model impanelments, depending on the specific requirements of the MapRL task.
2.3. Downstream Tasks
The downstream tasks consist of two components: Downstream Models and Fine-tuning Loss. Downstream models use the representation vectors generated by the encoders to perform specific tasks. Fine-tuning loss adjusts the encoder parameters to improve the adaptability of the representation vectors for these tasks. In contrast to pre-training tasks, which are closely tied to the encoders, downstream tasks are heavily influenced by the type of map entities. This subsection presents a taxonomy of downstream tasks.
Tasks for POIs. Common downstream tasks for POI entities fall into three categories:
  • POI Classification (POIC): This task predicts the category of a POI based on its representation. The downstream model typically uses an MLP network with a SoftMax classifier.
  • Next POI Prediction (NPP): This task converts a POI-based trajectory into a sequence of representation vectors using an encoder. The downstream model, often a deep sequential model like LSTM, predicts the next POI in the trajectory using the representation of the first half. The output is a multi-class classification, with each class representing a possible POI.
  • Trajectory User Link (TUL): This task predicts the user who generated a trajectory based on the trajectory’s representation sequence. The downstream model is a deep sequential model, with each output class corresponding to a user ID.
The fine-tuning loss for all three tasks is the cross-entropy loss.
Tasks for Road Segments. Common downstream tasks for road segment entities fall into three categories:
  • Average Speed Inference (ASI). This task leverages the representation vector to infer the average speed of a road segment. The average speed is derived from trajectory data and must be excluded from the input features of the representation encoders to prevent data leakage. Inferring average speed is a representative task for estimating unknown features of a road segment. The downstream model for this task is a linear regression function, and the fine-tuning loss function is Mean Square Error (MSE).
  • Travel Time Estimation (TTE). This task estimates the travel time from the origin to the destination along a trajectory composed of road segments. The downstream model is a deep sequential model. Its input is the representation sequence of the segment-based trajectory, and the output is the regressed travel time. The fine-tuning loss function is MSE.
  • Similarity Trajectory Search (STS). This task aims to find the most similar trajectory from a large database given a trajectory query. The downstream model first employs a deep sequential model to convert the representation sequences of all trajectories in the database into trajectory representation vectors. Then, it uses cosine similarity between these representations to identify the most similar trajectory to the query. The fine-tuning loss function is a contrastive learning loss, such as InfoNCE, which treats the ground-truth trajectory as the positive sample and others as negative samples. The ground-truth trajectory is typically generated by replacing a randomly selected sub-trajectory with another that shares the same start and endpoints (Jiang et al., 2023; Mao et al., 2022; Chang et al., 2023).
Tasks for Land Parcels. The downstream tasks for land parcel entities primarily focus on attribute classification, such as function and population, and mobility volume. These tasks are typically categorized into three types:
  • Land Parcel Classification (LPC): This task predicts specific labels for a land parcel, such as land usage or functional category, based on its representation vector. The downstream function is typically a MLP with a SoftMax classifier. The fine-tuning loss is cross-entropy.
  • Flow Inference (FI): This task predicts the incoming and outgoing population of a land parcel over a given period (e.g., one month) using its representation vector. The downstream function is usually a linear regression, with MSE as the loss.
  • Mobility Inference (MI): This task predicts the OD flow between pairs of land parcels. The input to the downstream function consists of the representations of two parcels, and the output is the predicted OD flow volume. The downstream function can be a neural network, with MSE as the fine-tuning loss.
There are two fine-tuning strategies for downstream tasks (Chang et al., 2023; Raffel et al., 2020). $i$) Downstream Fine-tuning: In this strategy, only the downstream models are fine-tuned, while the parameters of the MapRL encoders remain frozen. $ii$) End-to-End Fine-tuning: This approach tunes the parameters of both the downstream models and the MapRL encoders, enabling the entire model to be optimized jointly.
Figure 4. The overall pipeline for VecCity.
3. The Library – VecCity
Based on the method-based taxonomy of map entity representation learning models, we propose VecCity, an easy-to-use toolkit for MapRL model development. Figure 4 illustrates the framework of the VecCity library, comprising three functional modules aligned with the four key elements of MapRL models outlined in Section 2. $i$) Data Module: Corresponding to the map data element in Section 2.1, this module handles data preparation and processing. $ii$) Upstream Module: Aligned with the encoder and pre-training elements in Section 2.2, it is responsible for representation encoding and model pre-training. $iii$) Downstream Module: Linked to the downstream task element in Section 2.3, this module manages model fine-tuning and performance evaluation on downstream tasks.
3.1. Data Module and Atomic Files
In the Data Module, map entities and auxiliary data are represented in a unified file format, called Atomic Files (Wang et al., 2021). This format consists of three file types: .geo -– represents map entities, .traj -– stores trajectory data, .rel – captures relation networks.
.geo Files. A .geo file stores a list of map entities, such as POIs, road segments, and land parcels. Each line in the file represents an entity with four fields: geo_uid, geo_type, geo_location, and geo_features.
  • geo_uid. This field records the unique identifier of a map entity.
  • geo_type. This field indicates the entity type, which can be “point,” “polyline,” or “polygon,” corresponding to POIs, road segments, and land parcels, respectively.
  • geo_location. This field stores the geographical coordinates of the entity: a single longitude-latitude pair for points, a sequence of coordinates for polylines, and a closed sequence for polygons.
  • geo_features (optional). This field contains the entity’s features, with its format and length varying based on the data type and number of features.
.traj Files. A .traj file stores a set of map entity-based trajectories. Each line corresponds to a trajectory sample and contains four fields: traj_uid, sample_id, entity_id, and time_stamp.
  • traj_uid. Specifies the unique ID of the trajectory to which the sample belongs.
  • sample_order. Records the sequential index of the sample within the trajectory.
  • entity_id. Identifies the map entity corresponding to the sample, with its value matching the geo_uid in the .geo file.
  • time (optional). Stores the timestamp of the sample in Unix timestamp format.
.rel Files. The .rel file uses a sparse matrix format to represent the adjacency matrix of a relational network. Each line records a non-zero relationship between two map entities, containing the following fields: rel_uid, orig_geo_id, dest_geo_id, weight, feature, and timestamp.
  • rel_uid. Records the unique ID for the relationship between two map entities.
  • orig_geo_id and dest_geo_id. Specify the unique IDs of the origin and destination map entities involved in the relation, matching the geo_uid values in the .geo file.
  • weight. Stores the weight of the edge corresponding to the map entity relation.
  • feature (optional). Contains additional features for the edge. In some models, relation networks are represented as heterogeneous graphs, and this field can store the type of edges.
  • time (optional). Stores the timestamp for the relation, which is essential for dynamic graphs where edge weights and features may change over time.
Using the three types of atomic files, we can represent map entities and auxiliary data in a unified format. Original data often comes in various spatiotemporal formats, such as GeoJSON (Butler et al., 2016), Shapefile (Environmental Systems Research Institute, 1998), and NPZ (Harris et al., 2020), requiring multiple data-loading implementations and hindering reusability. Our VecCity library solves this issue by adopting unified atomic files, simplifying data processing.
In VecCity, we have pre-loaded map entities and auxiliary data from nine cities: New York (NY), Chicago (CHI), Tokyo (TYO), Singapore (SIN), Porto (PRT), San Francisco (SF), Beijing (BJ), Chengdu (CD), and Xi’an (XA), using public data sources. Map entities are extracted from OpenStreetMap (OSM), an open-source global map. POI check-in trajectories come from Foursquare datasets (Yang et al., 2016), while coordinate trajectories are collected from various studies (Jiang et al., 2023; did, 2022; Meghan O’Connell, 2015; Piorkowski and Grossglauser, 2009). The trajectory data preprocessing approach aligns with prior studies (Feng et al., 2018). Relation networks are OD flow networks constructed from these trajectory datasets. All datasets are processed, converted into atomic files, and integrated into the VecCity library. Details of the datasets are provided in Tab. 2. Users can directly apply these datasets in their model development. Additionally, we provide data conversion scripts on the VecCity GitHub repository, enabling users to convert their private data into the atomic file format.
Table 2. Pre-prepared datasets for the VecCity library.
City #POI #Segment #Parcel Check-in Traj Coor. Traj #OD
#Traj #User #Traj #User
NY 79,627 90,781 262 823,853 2,758 - - 28,189
CHI 28,141 47,669 77 279,183 921 - - 4,631
TYO 61,858 407,905 64 573,703 2,227 226,782 - -
SIN 75,329 35,084 332 827,828 3,489 11,170 - -
PRT 4,521 11,095 382 19,402 126 695,085 435 324
SF 15,674 27,274 194 171,873 361 500,516 405 24,716
BJ 81,181 40,306 11,208 - - 1,018,312 1,677 687,890
CD 17,301 6,195 1,306 - - 559,729 48,295 64,482
XA 19,108 5,269 1,056 - - 384,618 26,787 54,365
3.2. Upstream Module and Downstream Module
Upstream Module. Based on the taxonomy of MapRL encoders and pre-training tasks in Section 2.2, the VecCity library provides the upstream module through two key interfaces: encode() and pretraining_loss().
  • encode(): This interface implements the encoder for an MapRL model. It takes data extracted from atomic files as input and outputs representation vectors for map entities.
  • pretraining_loss(): This interface manages the implementation of pre-training tasks. It prepares samples through methods such as data masking, generating positive/negative samples, or data augmentation, and calculates the corresponding loss values.
In the VecCity framework, users only need to implement the interface functions in the upstream module to execute the pre-training phase of an MapRL model. Using these two interfaces, we have reproduced 21 mainstream MapRL models in VecCity, with 7 models each for POI, road segment, and land parcel representation learning. The names of the reproduced models are marked with “*” in Tab. 1.
Downstream Module. The downstream module provides two key interfaces: downstream_model() and finetuning_loss().
  • downstream_model(): This interface implements the downstream models used during the fine-tuning step. It typically utilizes a simple MLP or LSTM model. For classification tasks, the output layer is a SoftMax classifier, while for continuous prediction tasks, it is a linear regression layer.
  • finetuning_loss(): This interface defines the loss function for fine-tuning. Cross-entropy is used for classification tasks, MSE for regression tasks, and InfoNCE for contrastive learning tasks.
Using these two interfaces, we have implemented all nine downstream tasks listed in Section 2.3. Users can directly invoke these pre-set tasks to fine-tune their models.
Table 3. Performance comparison of POI MapRL methods on various datasets. The bold/underlined font means the best/the second-best result. "↑" means bigger is better and "↓" means smaller is better.
Dataset New York Tokyo Chicago Singapore San Francisco Overall
Task POIC NPP POIC NPP POIC NPP POIC NPP POIC NPP Pre Ave Rank Pre Rank
Metric ACC@1↑ F1↑ ACC@1↑ ACC@5↑ ACC@1↑ F1↑ ACC@1↑ ACC@5↑ ACC@1↑ F1↑ ACC@1↑ ACC@5↑ ACC@1↑ F1↑ ACC@1↑ ACC@5↑ ACC@1↑ F1↑ ACC@1↑ ACC@5↑
SkipG 0.071 0.012 0.126 0.309 0.334 0.042 0.164 0.333 0.077 0.007 0.050 0.161 0.027 0.003 0.030 0.074 0.011 0.003 0.014 0.053 5.3 7
Tale 0.057 0.009 0.152 0.341 0.212 0.037 0.145 0.297 0.096 0.009 0.094 0.230 0.029 0.005 0.043 0.101 0.009 0.001 0.007 0.046 4.8 6
Teaser 0.079 0.014 0.131 0.329 0.354 0.048 0.137 0.241 0.077 0.005 0.029 0.137 0.065 0.018 0.033 0.078 0.037 0.012 0.014 0.058 4.4 5
Hier 0.052 0.008 0.121 0.285 0.278 0.032 0.162 0.335 0.086 0.011 0.036 0.161 0.041 0.006 0.079 0.163 0.039 0.007 0.012 0.108 4.2 4
P2Vec 0.074 0.013 0.116 0.154 0.313 0.037 0.154 0.319 0.111 0.012 0.143 0.393 0.068 0.014 0.075 0.168 0.033 0.011 0.019 0.046 3.7 3
CACSR 0.028 0.004 0.135 0.331 0.232 0.010 0.159 0.325 0.104 0.017 0.146 0.394 0.030 0.005 0.103 0.126 0.018 0.004 0.077 0.158 3.5 2
CTLE 0.048 0.006 0.154 0.336 0.235 0.014 0.168 0.338 0.119 0.019 0.187 0.398 0.083 0.008 0.107 0.224 0.036 0.006 0.060 0.123 2.0 1
Task TUL Efficiency TUL Efficiency TUL Efficiency TUL Efficiency TUL Efficiency Eff Ave Rank Eff Rank
Metric ACC@1↑ F1↑ Param↓ Speed↓ ACC@1↑ F1↑ Param↓ Speed↓ ACC@1↑ F1↑ Param↓ Speed↓ ACC@1↑ F1↑ Param↓ Speed↓ ACC@1↑ F1↑ Param↓ Speed↓
SkipG 0.625 0.317 2.632 2 0.452 0.243 2.013 3 0.314 0.085 0.937 1 0.158 0.047 2.513 3 0.123 0.138 0.414 1 1.0 1
Tale 0.699 0.429 8.611 5 0.395 0.219 8.111 12 0.609 0.360 2.910 2 0.300 0.164 9.919 13 0.329 0.151 1.107 1 4.4 6
Teaser 0.608 0.307 3.012 3 0.464 0.255 2.318 5 0.312 0.193 1.064 1 0.173 0.152 2.986 5 0.556 0.184 0.464 1 1.8 2
Hier 0.635 0.347 8.253 5 0.493 0.251 7.631 8 0.521 0.289 6.550 2 0.136 0.142 8.133 9 0.557 0.242 6.025 1 4.2 4
P2Vec 0.629 0.323 7.142 7 0.451 0.243 2.735 12 0.647 0.413 1.410 2 0.386 0.169 3.346 11 0.517 0.256 1.496 1 3.6 3
CACSR 0.645 0.431 9.921 43 0.502 0.249 8.301 114 0.659 0.462 5.439 21 0.361 0.156 9.715 68 0.611 0.295 4.059 7 6.7 7
CTLE 0.698 0.444 3.436 8 0.507 0.294 2.814 13 0.772 0.628 1.734 4 0.485 0.243 3.316 12 0.628 0.323 1.209 1 4.2 4
3.3. Evaluation & Config Modules
The evaluation module provides an evaluation() interface to measure the performance of MapRL models on downstream tasks. the VecCity library offers a range of metrics tailored to different tasks. Specifically, it includes MAE, MSE, RMSE, R2, and MAPE for regression tasks; Accuracy, Precision@k, Recall@k, F1@k, and Mean Rank (MR) for classification tasks; and Accuracy@k and MR for trajectory similarity search tasks.
The config module provides a configuration file as an interface for users to arrange the “pre-training $\rightarrow$ fine-tuning $\rightarrow$ evaluation” pipeline. To train and test a model using VecCity, users simply specify the required pre-built or custom models and functions for each step in the pipeline within the configuration file. The library then automatically generates the necessary script to run experiments based on this configuration. Users only need to execute the script, and VecCity will invoke the interface functions of each module to complete the experiment.
4. Performance Benchmark
In this section, we present performance comparison experiments for the pre-built MapRL models in VecCity. These experiments serve three purposes: $i$) Demonstrating Usability and Power: To show that VecCity is a user-friendly and powerful toolkit for MapRL model development and evaluation. $ii$) Establishing Performance Benchmarks: To provide standard performance benchmarks for mainstream MapRL models using the datasets pre-loaded in VecCity. $iii$) Analyzing Component Contributions: To explore the contributions of different components in existing MapRL models, offering insights for designing more effective and robust models.
4.1. Datasets and Experimental Setups
4.1.1. Datasets
The dataset selection for our experiments follows these guiding principles: for each type of MapRL model, we incorporate all datasets that meet the task requirements. Specifically, for POI-oriented MapRL models, we use datasets from New York City, Chicago, Tokyo, Singapore, and San Francisco. For road-segment-oriented models, the datasets include Beijing, Xi’an, Chengdu, Porto, and San Francisco. For land-parcel-oriented models, we adopt datasets from Beijing, Xi’an, Chengdu, Porto, and San Francisco. Each dataset is divided into training, validation, and test sets following a 6:2:2 ratio, ensuring balanced and reliable performance evaluation. For each downstream task, we apply end-to-end fine-tuning to achieve optimal performance.
4.1.2. Experimental Settings
In our experiments, the dimension of map entity representation vectors across all models is set to 128. For other model configurations, such as the number of layers and the dimensions of intermediate variables, we adopt the default settings from the original papers as initial values. We then use a trial-and-error approach to tune the parameters on each dataset, searching for the best performance configurations. All experiments are executed on an Ubuntu 20.04 system equipped with NVIDIA GeForce 3090 GPUs. The default batch size during model training is set to 64. In cases of out-of-memory (OOM) issues, the batch size is halved iteratively, with a minimum batch size of 8.
Table 4. Performance comparison of road segment MapRL methods on various datasets.
Dataset Beijing Chengdu Xi’an Porto San Francisco Overall
Task ASI TTE ASI TTE ASI TTE ASI TTE ASI TTE Per Ave Rank Per Rank
Metric MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓ MAE↓ RMSE↓
RN2Vec 3.566 5.744 656.6 3050 6.490 13.90 105.9 154.8 6.237 12.72 130.38 190.5 4.380 7.896 97.8 144.5 2.978 5.797 289.0 1479 6.0 7
HRNR 3.157 5.508 625.1 3066 6.364 13.66 100.26 149.78 5.672 11.11 129.24 191.1 4.357 7.867 104.5 150.2 2.510 5.374 287.8 1445 5.8 6
SARN 2.798 5.479 566.8 2578 6.025 13.41 95.64 148.3 5.315 10.60 117.0 178.7 4.286 7.779 99.5 145.8 2.481 5.303 279.0 1502 4.1 5
Toast 2.837 5.512 513.0 2499 6.377 13.62 90.7 145.4 5.619 10.96 116.3 184.9 4.354 7.883 96.4 144.6 2.553 5.295 246.4 1270 3.6 4
HRoad 2.726 5.411 610.3 2527 5.970 13.22 96.24 146.6 5.537 10.81 116.6 187.4 4.212 7.693 97.2 144.3 2.476 5.250 287.6 1546 3.5 3
START 2.738 5.445 509.4 2489 6.066 13.47 92.3 145.9 5.503 10.92 116.4 183.4 4.210 7.792 95.8 145.7 2.450 5.333 261.8 1431 2.6 2
JCLRNT 2.771 5.466 475.9 2442 6.354 13.27 90.2 142.7 5.660 11.02 115.9 183.5 4.267 7.699 94.6 143.4 2.534 5.368 264.5 1416 2.4 1
Task STS Efficiency STS Efficiency STS Efficiency STS Efficiency STS Efficiency Eff Ave Rank Eff Rank
Metric ACC@3↑ MR↓ Param↓ Speed↓ ACC@3↑ MR↓ Param↓ Speed↓ ACC@3↑ MR↓ Param↓ Speed↓ ACC@3↑ MR↓ Param↓ Speed↓ ACC@3↑ MR↓ Param↓ Speed↓
RN2Vec 0.789 10.53 5.192 331 0.631 45.32 0.708 19 0.695 11.208 0.826 16 0.893 8.296 1.453 3 0.976 6.179 3.524 35 2.7 2
HRNR 0.809 9.915 5.699 281 0.698 37.12 1.188 5 0.709 10.580 1.321 4 0.807 11.170 1.948 80 0.915 6.988 4.066 23 3.6 4
SARN 0.892 9.176 0.570 30 0.563 40.31 0.549 5 0.686 7.384 0.551 5 0.883 8.753 0.551 10 0.963 6.879 0.563 19 1.2 1
Toast 0.828 9.755 10.01 495 0.678 18.64 1.840 192 0.749 5.874 1.566 132 0.895 9.031 3.045 252 0.978 5.934 7.021 253 6.0 6
HRoad 0.947 8.491 5.398 66 0.583 32.05 0.913 6 0.736 7.122 1.031 5 0.895 10.152 1.660 11 0.947 8.491 3.728 44 3.3 3
START 0.948 7.867 6.238 1719 0.713 15.59 2.139 582 0.755 5.411 2.002 312 0.905 6.702 2.745 735 0.934 6.828 4.743 783 6.7 7
JCLRNT 0.879 9.034 5.114 434 0.724 14.34 1.049 287 0.797 4.746 0.912 150 0.927 4.778 1.649 345 0.957 6.284 3.630 198 4.3 5
4.2. Overall Experimental Results
4.2.1. Experimental Results for POI MapRL
Table 3 presents the performance of MapRL models on downstream tasks, including POI Classification (POIC), Next POI Prediction (NPP), and Trajectory User Link (TUL). Each task is evaluated using two types of metrics. Additionally, model efficiency is assessed with two indicators: the number of parameters (Param) in millions and the time cost per epoch during pre-training (Speed) in seconds. The table also reports average performance rankings across tasks (Pre Ave Rank) and the ranking based on the models’ average performance (Pre Rank). Similarly, efficiency rankings are given by Eff Ave Rank and Eff Rank. From Tab. 3, we have the following observations.
1) Importance of Sequence-based Encoders: Models combining token-based and sequence-based encoders – such as CTLE, CACSR, and Hier – consistently outperform those relying solely on token-based encoders, highlighting the value of trajectory auxiliary data. Among these, models using Augmented Trajectory Contrastive Learning (ATrCL) and Masked Trajectory Recovery (MTR) tasks (e.g., , CTLE and CACSR) achieve better performance than those focused on Trajectory Prediction (TrajP) tasks (e.g., , Hier). This is likely because ATrCL and MTR capture long-range temporal dependencies across entire trajectories, while TrajP focuses only on consecutive POI dependencies.
2) Impact of Relation Networks: For token-based models, those incorporating relation networks – such as Teaser and P2Vec – outperform models without them, such as SkipGram and Tale, underscoring the value of relation network auxiliary data. However, for models with both token-based and sequence-based encoders, the inclusion of relation networks offers little additional benefit. This is likely due to the sparsity of POI trajectories, where the number of trajectories between two POIs is often dictated by geographical proximity. As a result, the information provided by relation networks overlaps significantly with that from trajectory data. Once the dependency within the trajectory data has been fully exploited, the relational network contributes limited additional information.
3) TokRI vs. TRCL Tasks: Token-based models using the Token Relation Inference (TokRI) task, such as Tale and P2Vec, relatively outperform those employing the Token Relation Contrastive Learning (TRCL) task, such as SkipGram and Teaser. With sufficient data, supervised learning via TokRI effectively leverages labeled data to generate more discriminative feature representations. In contrast, the indirect nature of contrastive learning in TRCL offers less direct utilization of labeled information, resulting in comparatively lower model performance.
4) Efficiency of Token-based Models: Token-based models exhibit lower computational complexity, making them more efficient. This makes them especially suitable for application scenarios where computational resources are limited, and performance requirements can be moderately compromised.
4.2.2. Experimental Results for Road Segment MapRL
Table 4 shows a performance comparison of road segment MapRL models across three downstream tasks: Average Speed Inference (ASI), Travel Time Estimation (TTE), and Similarity Trajectory Search (STS). The following key observations can be drawn from the results:
1) Full Pipeline Models Perform Best: Models utilizing the complete token-graph-sequence pipeline, such as JCLRNT and START, achieve the highest performance compared to models with partial pipelines. Additionally, token-based + graph-based and token-based + sequence-based models – such as HRNR, SARN, Toast, and HRoad – outperform purely token-based models like RN2Vec, highlighting the value of integrating trajectory and relation network information. Graph-based and sequence-based encoders capture essential patterns that improve road segment representations.
2) Multiple Pre-training Tasks Improve Performance: Models with multiple types of pre-training tasks, i.e., Toast, HRoad, START, and JCLRNT, outperform those with single-task pre-training, i.e., HRNR and SARN. Notably, among token-based + graph-based models, HRoad, which integrates both NFI and GAu tasks, surpasses HRNR (GAu only) and SARN (AGCL only). Using diverse pre-training tasks enables the model to extract information from multiple perspectives, enhancing its ability to learn more comprehensive feature representations. Moreover, in the full-pipeline models, JCLRNT combines a sequence-oriented task (Augmented Trajectory Contrastive Learning) with a graph-oriented task (Augmented Graph Contrastive Learning) and outperforms START, which uses two sequence-oriented tasks (MTR and ATrCL). This suggests that even in models with multiple tasks, greater task heterogeneity can further enhance performance.
3) Efficiency Trade-offs: Simpler models with fewer modules, such as SARN and RN2Vec, are significantly more efficient. Models without sequence encoders, such as HRNR and HRoad, also exhibit better efficiency since incorporating trajectory data demands considerable memory and time. Improving the efficiency of trajectory-based models remains a key challenge for future research.
Table 5. Performance comparison of land parcel MapRL methods on various datasets.
Dataset Beijing Chengdu Xi’an Porto San Francisco Overall
Task LPC FI LPC FI LPC FI LPC FI LPC FI Per Ave Rank Per Rank
Metric ACC@1↑ F1↑ MAE↓ RMSE↓ ACC@1↑ F1↑ MAE↓ RMSE↓ ACC@1↑ F1↑ MAE↓ RMSE↓ ACC@1↑ F1↑ MAE↓ RMSE↓ ACC@1↑ F1↑ MAE↓ RMSE↓
ZEMob 0.380 0.110 30.90 55.19 0.339 0.115 154.9 441.2 0.316 0.138 133.7 326.2 0.298 0.204 527.5 595.1 0.820 0.477 509.1 910.5 5.2 7
GMEL 0.412 0.571 27.68 58.91 0.338 0.109 154.8 441.1 0.330 0.359 133.1 324.4 0.267 0.335 525.1 591.9 0.620 0.550 514.0 916.7 4.5 6
MGFN 0.567 0.601 30.09 54.45 0.345 0.135 154.6 448.2 0.330 0.162 132.7 324.9 0.304 0.246 531.3 596.9 0.835 0.533 605.3 1011 4.4 5
HDGE 0.409 0.229 29.26 54.82 0.365 0.157 159.6 402.1 0.473 0.242 134.1 334.5 0.327 0.263 525.2 576.9 0.820 0.450 511.8 926.4 4.3 4
MVURE 0.479 0.267 28.13 51.37 0.485 0.279 161.5 450.2 0.545 0.235 133.7 326.0 0.316 0.294 524.8 592.4 0.820 0.450 511.2 913.1 4.1 3
ReMVC 0.722 0.686 27.81 52.66 0.610 0.542 147.7 430.9 0.567 0.412 131.5 324.6 0.404 0.357 527.3 594.8 0.820 0.450 444.7 852.3 3.0 2
HREP 0.524 0.405 22.91 45.85 0.528 0.445 124.0 309.6 0.611 0.494 127.9 312.3 0.386 0.358 498.0 574.5 0.836 0.538 438.4 799.3 2.2 1
Task MI Efficiency MI Efficiency MI Efficiency MI Efficiency MI Efficiency Eff Ave Rank Eff Rank
Metric MAE↓ RMSE↓ Param↓ Speed↓ MAE↓ RMSE↓ Param↓ Speed↓ MAE↓ RMSE↓ Param↓ Speed↓ MAE↓ RMSE↓ Param↓ Speed↓ MAE↓ RMSE↓ Param↓ Speed↓
ZEMob 0.379 0.771 45.10 40 3.599 29.66 7.699 19 3.776 17.63 5.670 1 44.85 72.30 1.047 1 5.837 23.48 2.335 1 4.0 3
GMEL 0.418 0.712 0.118 713 4.599 29.66 0.118 193 3.668 17.60 0.118 189 41.08 57.41 0.118 161 7.377 22.29 0.118 2 4.0 3
MGFN 0.283 0.694 1156 839 5.176 29.23 485.7 6 3.889 17.23 318 2 39.37 56.16 53.67 2 6.846 22.22 42.19 1 5.4 7
HDGE 0.389 0.759 12.91 1799 5.189 29.00 1.505 601 3.863 16.85 1.217 75 38.69 54.20 0.440 117 7.397 21.94 0.223 3 5.1 6
MVURE 0.281 0.685 0.097 12 5.200 29.37 0.097 8 3.941 17.41 0.097 1 38.63 62.21 0.097 1 7.196 22.73 0.097 1 2.1 1
ReMVC 0.276 0.667 1.450 873 5.230 29.42 0.181 582 3.986 17.40 0.149 39 40.34 56.66 0.069 47 7.007 22.60 0.050 7 4.1 5
HREP 0.318 0.677 0.182 9 5.270 28.33 0.182 7 3.949 17.05 0.182 1 37.42 57.62 0.182 2 6.754 21.71 0.182 1 3.3 2
4.2.3. Experimental Results for Land Parcel MapRL
Table 5 presents the performance comparison of land parcel MapRL models on the downstream tasks of Land Parcel Classification (LPC), Flow Inference (FI), and Mobility Inference (MI). The following observations can be made:
1) Incorporating POI Data Enhances Performance: Models that utilize POI data, such as MVURE, ReMVC, and HREP (see the “Map Entities” column in Table 1), outperform those that do not. These models treat POI categories within a land parcel as features, enriching the representation of the parcel’s functional characteristics. This approach provides valuable contextual information, improving the overall performance of the models.
2) Limitations of Current Encoder Structures: Most existing land parcel MapRL models adopt a token-based + graph-based encoder structure, focusing on extracting information from relation networks. However, these models overlook the sequential dependencies within trajectories, limiting their effectiveness. Future research should explore advanced methods to capture trajectory sequence information relevant to land parcels for further improvements.
3) Need for More Diverse Pre-training Tasks: Current models mainly rely on the Graph Autoencoder (GAu) task for pre-training. However, experimental results show that tasks such as Augmented Token Contrastive Learning (AToCL) and Neighborhood Contrastive Learning (NCL) yield better performance. Developing models that incorporate a broader range of pre-training tasks is a promising direction for enhancing land parcel MapRL models.
4.2.4. Summary for Overall Experiments
The overall experimental results highlight the following key findings across POI, road segment, and land parcel MapRL models.
1) Importance of Combined Encoders: Models employing a combination of token-based, graph-based, and sequence-based encoders consistently outperform those relying on single-type encoders. This demonstrates the value of integrating multiple data sources – such as trajectories and relation networks – to capture complex spatial, temporal, and relational patterns.
2) Effectiveness of Diverse Pre-training Tasks: Models that leverage multiple pre-training tasks usually perform better than those limited to a single task. This suggests that diverse pre-training strategies enhance the model’s ability to learn comprehensive feature representations from different perspectives.
3) Role of Auxiliary Data: Incorporating POI, trajectory, and relation network data improves model performance across tasks. However, models that fully exploit trajectory dependencies achieve diminishing gains from relation networks, indicating some redundancy between these data sources.
4) Trade-offs Between Performance and Efficiency: While complex models with multiple encoders and tasks offer higher accuracy, they require more computational resources. Simpler token-based models remain attractive for applications where efficiency is critical, despite offering slightly lower performance.
4.3. Performance over Small Datasets
In the overall experiments, both the pre-training and fine-tuning steps are executed on the full training dataset. However, in practical applications, MapRL models are often trained on abundant data while fine-tuned using limited labeled samples for downstream tasks. To assess the robustness of mainstream models under such conditions, we conduct a small dataset experiment.
In this experiment, we reduce the fine-tuning data size from 50% to 10% of the original training set and compare the result of pre-built models in VecCity. POI-oriented models are evaluated on the Tokyo dataset, while road-segment and land-parcel-oriented models are tested on the Xi’an dataset. Figure 5 presents the evaluation result: Fig. 5 (a) to (c) show results for POI-oriented models. Fig. 5 (d) to (f) display results for road-segment-oriented models. Fig. 5 (g) to (i) provide results for land-parcel-oriented models. From these figures, we make following observations:
Figure 5. Results from varying fine-tuning data size.
1) Stability in Attribute Inference Tasks: MapRL models show stable performance across varying training data sizes in attribute inference tasks, such as POI Classification (POIC) in Fig. 5 (a), Average Speed Inference (ASI) in Fig. 5 (d), and Land Parcel Classification (LPC) in Fig. 5 (g). This stability arises because these tasks focus on inferring entity attributes, which are effectively captured during pre-training. As a result, the fine-tuning phase requires only a small amount of labeled data to maintain good performance.
2) Fluctuations in Trajectory-Related Tasks: Models exhibit significant performance fluctuations in trajectory-related tasks, such as Next POI Prediction (NPP) in Fig. 5 (b), Trajectory User Link (TUL) in Fig. 5 (c), and Travel Time Estimation (TTE) in Fig. 5 (e). These tasks rely on sequence-level predictions, while encoders and pre-training tasks focus on generating single-vector representations for individual entities. This mismatch makes trajectory-related tasks more sensitive to the size of the fine-tuning dataset, as limited labels reduce the model’s ability to learn sequential dependencies effectively. For Similarity Trajectory Search (STS) in Fig. 5 (f), although it is a trajectory-related task, it involves similarity comparison rather than prediction so is less affected by this mismatch.
3) Fluctuations in Flow and Mobility Inference Tasks: Models also show significant performance fluctuations in Flow Inference (FI) in Fig. 5 (h) and Mobility Inference (MI) in Fig. 5 (i), even though these tasks are not directly related to trajectory data. This is because the labels for these tasks reflect statistical data on human mobility flows over specific periods, which are highly time-heterogeneous, with label distributions varying significantly across periods. As a result, more labeled data is required to capture these complex temporal patterns accurately.
Overall, existing MapRL models exhibit varying sensitivity to the size of labeled data across different tasks. For tasks less influenced by temporal information, such as attribute inference (ASI, POIC, LPC) and Similarity Trajectory Search (STS), models maintain stable performance even with limited training data. However, for tasks with time-dependent labels, such as trajectory-related tasks (NPP, TUL, TTE) and flow/mobility inference tasks (FI, MI), larger amounts of labeled data are necessary to effectively capture temporal dynamics.
Table 6. Results for CTLE variants on the Tokyo dataset.
Methods CTLE CTLECTLE& CTLE* CTLE\$
Encoder Core Arc TE TE TE TE LSTM
Pretraining Task TrajP
MTR
ATrCL
POIC ACC@1↑ 0.048 0.032 0.041 0.050 0.043
F1↑ 0.006 0.002 0.005 0.007 0.006
NPP ACC@1 0.154 0.089 0.153 0.165 0.058
ACC@5↑ 0.336 0.193 0.332 0.352 0.147
TUL ACC@1 0.698 0.401 0.687 0.699 0.388
F1↑ 0.444 0.160 0.437 0.448 0.104
Per Rank 2 5 3 1 4
4.4. Comparing of Modified Models
In this section, we conduct experiments to evaluate the influence of pre-training tasks and encoder structures on the performance of downstream tasks. We modify three representative MapRL models – CLTE, JCLRNT, and HREP – which achieve the highest average performance in POI, road segment, and land parcel representation learning, respectively. Owing to VecCity’s modular design, specific pre-training tasks can be easily removed from models, and new pre-training tasks can be seamlessly integrated.
The experimental results are shown in Tab. 68. In the figures, the black “✓” and “✗” indicate whether the corresponding pre-training tasks were originally included or excluded when the model was first proposed. In contrast, the red “” and “” indicate newly added or removed pre-training tasks. Additionally, we evaluate the impact of different backbone models by modifying the core architecture of encoders (highlighted in red in the “Encoder Core Arc” row). In the last row, we present the performance ranking of each variant for easier comparison and conclusion.
Comparison of CTLE Variants. Among POI MapRL models, CTLE achieves the best performance. In Tab. 6, we experiment with the following model variants for CTLE. (a) CTLE: Maintains the original model architecture. (b) CTLE%: Removes the Masked Trajectory Recovery (MTR) pre-training task. (c) CTLE&: Removes the Trajectory Prediction (TrajP) pre-training task. (d) CTLE*: Integrates a new Augmented Trajectory Contrastive Learning (ATrCL) pre-training task, using the ATrCL task proposed by CACSR, which achieved the second-best performance in the overall POI experiments. (e) CTLE\$: Replaces the core architecture of CTLE (a Transformer encoder) with an LSTM layer.
The downstream task results are presented in Tab. 6. From these experiments, we draw the following observations: $i$) New Task Enhances Performance: The CTLE* variant achieves the best performance across all downstream tasks, indicating that integrating the new ATrCL task significantly improves the model’s capability. $ii$) MTR Provides Richer Representations: The CTLE& variant outperforms CTLE%, suggesting that the MTR task offers richer representation capabilities than the TrajP task. $iii$) Transformer Encoder is Essential for Trajectory Tasks: The performance of CTLE\$ drops by 59% in the Next POI Prediction (NPP) task and 60% in the Trajectory User Link (TUL) task, showing that the Transformer Encoder (TE) plays a crucial role in trajectory-related tasks.
Table 7. Results for JCLRNT variants on the Xi’an dataset.
Methods JCLRNT JCLRNTJCLRNT& JCLRNT* JCLRNT\$
Encoder Core Arc TE TE TE TE LSTM
GAT GAT GAT GAT GAT
Pretraining Task TokRI
AGCL
ATrCL
ASI MAE↓ 5.660 5.708 5.744 5.601 5.702
RMSE↓ 11.02 11.06 11.11 10.90 11.00
TTE MAE↓ 115.9 118.3 116.5 115.3 124.5
RMSE↓ 183.5 185.5 183.8 182.3 189.8
STS ACC@3↑ 0.797 0.673 0.702 0.809 0.782
MR↓ 4.746 12.30 5.828 3.176 5.325
Per Rank 2 5 4 1 3
Comparison of JCLRNT Variants. Among road segment MapRL models, JCLRNT achieves the best performance. As shown in Tab. 7, we experiment with the following JCLRNT variants. (a) JCLRNT: Maintains the original model architecture. (b) JCLRNT%: Removes the ATrCL pre-training task. (c) JCLRNT&: Removes the Augmented Graph Contrastive Learning (AGCL) pre-training task. (d) JCLRNT*: Adds a new Token Relation Inference (TokRI) pre-training task, using road type classification, as implemented in HRoad. (e) JCLRNT\$: Replaces the Transformer with an LSTM as the sequence encoder.
The performance results on downstream tasks are presented in Tab. 7. From the experimental results, we draw the following observations: $i$) Comprehensive Pre-training Boosts Performance: JCLRNT* achieves the best performance across all downstream tasks by incorporating three pre-training tasks—TokRI, AGCL, and ATrCL—each aligned with the token-based, graph-based, and sequence-based encoders, respectively. This suggests that each encoder type benefits from having a tailored pre-training task. $ii$) ATrCL Provides Richer Representations: JCLRNTunderperforms compared to JCLRNT& in trajectory-related tasks like Travel Time Estimation (TTE) and Similarity Trajectory Search (STS). This indicates that ATrCL, a sequence-based pre-training task, offers richer representation capabilities than AGCL for these tasks. Even with a sequence encoder, the absence of ATrCL limits the model’s ability to capture dependency effectively, impacting performance. $iii$) Minor Impact of Replacing the Transformer Encoder: JCLRNT\$ shows relatively minor performance degradation when the Transformer Encoder is replaced by an LSTM. This suggests that the introduction of the graph encoder in JCLRNT helps extract relational information from network data, reducing the model’s reliance on the Transformer.
Table 8. Results from HREP variants on the Xi’an dataset.
Methods HREP HREPHREP& HREP* HREP\$
Encoder Core Arc HeteGCN HeteGCN HeteGCN HeteGCN GCN
Pretraining Task NFI
GAu
NCL
LPC ACC@1↑ 0.611 0.566 0.413 0.663 0.505
F1↑ 0.494 0.436 0.273 0.495 0.340
FI MAE↓ 127.9 128.1 133.4 119.7 146.6
RMSE↓ 312.3 320.8 323.7 269.8 325.3
MI MAE↓ 3.949 3.993 4.371 3.914 4.294
RMSE↓ 17.05 18.58 17.69 15.62 16.55
Per Rank 2 3 5 1 4
Comparison of HREP Variants. Among land parcel MapRL models, HREP achieves the best performance. As shown in Tab. 8, we experiment with the following HREP variants. (a) HREP: Maintains the original model architecture. (b) HREP%: Removes the Graph Autoencoder (GAu) pre-training task. (c) HREP&: Removes the Neighborhood Contrastive Learning (NCL) pre-training task. (d) HREP: Adds a new Node Feature Inference (NFI) pre-training task, incorporating human mobility inference, as implemented in MGFN. (e) HREP\$: Replaces the heterogeneous GCN (HeteGCN) with a standard GCN as the graph encoder.
The performance results on downstream tasks are presented in Tab. 8. From the experiments, we draw the following observations. $i$) NFI Enhances Flow Inference Performance: HREP* achieves the best results across all downstream tasks, including a 10% improvement in the Flow Inference (FI) task, indicating that the NFI pre-training task significantly enhances the model’s ability to predict flow-related information. $ii$) GAu Provides Richer Representations: HREPoutperforms HREP&, suggesting that the GAu pre-training task is crucial for capturing relationships between entities and improving downstream task performance. $iii$) HeteGCN Captures Diverse Entity Relationships: HREP\$ experiences a significant performance drop, especially in the Land Parcel Classification (LPC) task, when the HeteGCN is replaced with a GCN. This indicates that HeteGCN better captures the diverse relationships between different types of entities, making it essential for land parcel representation learning.
Summary for Variant Experiments. From the above experiments, we derive the following conclusions.
1) More Types of Pre-training Tasks Lead to Better Performance: Across all experiments, the best-performing models consistently incorporate the most diverse types of encoders and pre-training tasks. This demonstrates that integrating a wider range of pre-training tasks enhances the model’s ability.
2) Suitable Pre-training Tasks Provide Greater Benefits: Pre-training tasks tailored to specific downstream tasks significantly enhance performance. For instance, the Augmented Trajectory Contrastive Learning (ATrCL) task improves performance in trajectory-relatd tasks, such as Travel Time Estimation (TTE) and Similarity Trajectory Search (STS), while the Node Feature Inference (NFI) task greatly benefits Flow Inference (FI) tasks. This highlights the importance of aligning pre-training tasks with the specific requirements of downstream tasks to maximize model effectiveness.
3) Encoder Architecture Matters for Task Performance: The choice of encoder architecture has a significant effect. For instance, replacing the Transformer Encoder with an LSTM in CTLE and JCLRNT led to performance drops, especially in trajectory-related tasks. Similarly, HeteGCN outperforms a standard GCN in HREP by capturing more diverse entity relationships.
5. related work
Review for MapRL. Pervious reviews (Atluri et al., 2018; Wang et al., 2022; Jin et al., 2023) discuss the applications of deep learning in urban data mining, providing a detailed overview of data, tasks, and deep learning models, particularly those that integrate sequence and graph encoders. However, their focus is on end-to-end models, while our work concentrates on pre-trained representation learning methods. Other studies (Zheng, 2015; Chen et al., 2024d) focus on data mining methods for trajectory data but do not cover pre-trained representation learning. Survey (Chen et al., 2024a) introduces a type-based taxonomy for MapRL where models are systematically reviewed based on the types of map entities they target. In contrast, our work does not classify methods by entity type but instead focuses on summarizing general techniques applicable to all map entities, i.e., a method-based taxonomy. This broader perspective helps uncover common design principles underlying MapRL methods.
Benchmark for MapRL. In recent years, the growing interest in MapRL has created a strong demand for open benchmarks to comprehensively analyze baseline models. Open-source benchmarking is valuable not only for advancing research but also for enabling practitioners to easily evaluate models and apply them to open datasets. While significant progress has been made in benchmarking for multivariate time series prediction (Wang et al., 2021; Shao et al., 2023; Qiu et al., 2024; Wang et al., 2024), computer vision (Kai Chen, 2019; Deng et al., 2009; Lin et al., 2014), and neural language processing (Pranav Rajpurkar, 2018; Conneau and Kiela, 2018; Reddy et al., 2019; Wang et al., 2019), benchmarking in the MapRL field remains largely unexplored. To the best of our knowledge, VecCity is the first benchmark specifically for MapRL, allowing researchers to conduct detailed comparative experiments and develop new models.
6. Conclusion
To overcome the challenges of fragmentation and the lack of standardized benchmarks in MapRL, we introduced a novel taxonomy organized by functional modules rather than entity types. Building on this taxonomy, we developed VecCity, a modular library offering interfaces for encoding, pre-training, fine-tuning, and evaluation. Using VecCity, we established the first standardized benchmarks by reproducing 21 mainstream models and integrating datasets from nine cities. Moreover, we also conduct an in-depth evaluation of mainstream MapRL models. Our comprehensive experiments highlight the impact of pre-training tasks and encoder architectures, demonstrating the advantages of combining multiple components. VecCity provides a unified framework that promotes reusability, streamlines experimentation, and advances research in MapRL. Although VecCity currently supports most mainstream MapRL models, we plan to expand it further to cover a wider range of pre-trained spatiotemporal data representation learning algorithms, such as trajectory representation learning.